Coalite Meeting Update -2/3/17

 Meeting minutes / notes from 2 March 2017 – @ 6:30pm Arkwright Centre

Progress to date / Planned work and Issues. Points raised –

  • All above ground structures removed except Guards House and Pump House (pipe work supports existing equipment currently used on site), so will be the last buildings standing.
  • Buildings containing dust / asbestos have all been removed from site (2 bins of hazardous waste from this have been removed).
  • Recorded site activity – visual aid document (before and after) was handed out during the meeting and discussed. Yellow areas for access and the Red areas those were concrete, earth over ground structures.
  • Soils treatment areacontinues to “work”, however residual odours continue to be emitted.
  • Some concrete slabs will be crushed and reused in the lower levels of construction work / bases for buildings.
  • Odour sheeting / plastic placed over to limit odour, also spray polymer foam, odour control as well.
  • No sheets have been removed yet, this will have massive impact on odour peak when they are lifted.
  • The PID detectors are not selective to VOCs, they only know that the odours are from the chlorinated phenols family, some known as the dirty dozen of persistent pollutants.
  • Standing water (colour of pink) continues to be a progressive problem, especially following wet weather of recent. Waste water is continually pumped from site into tankers and is sent to West Bromwich for treatment.
  • The Environmental monitors (5 of) have worked and have been recording data. Storm Doris knocked one over and this has been placed back into the ground closer to the road on buttermilk lane.
  • There have been few complaints over the last few weeks. (phone calls EA, local authorities, face book etc), although 13th 14th 15th Feb, there were a few.
  • Complaints Talley: 24 hour number (49 complaints since start) and Env Health Dept (35 in total).
  • The monitors are; 2 on Coalite site, 1 at the Allotments, 1 at the Bolsover Industrial park and 1 near the River.
  • The monitor at the allotments was triggered. This was investigated and later found to be effected by a nearby garden fire.
  • Other monitors have detected consistent low levels; controlled VOCs were intentionally placed underneath one to make sure that it was still working / sampling data. It was, a sharp peak in levels were recorded at that point. Environmental Health dept. was told.
  • Bolsover land Ltd /DDC ensure that all activities are safe. They rather not be hasty in works undertaken, there is plenty of leeway for this.
  • Completion date perhaps around 21st April, but this will be reviewed.
  • Challenges: The area of soils with the odour potential is larger in area and higher than expected. The area of contamination is higher than site investigation initially evaluated.
  • The water treatment plant has not been touched yet. This may be one of those other challenges.
  • Although continued concerns over site and contamination issues, the site aesthetically looks much better with Tanks and buildings removed. More positive comments past.
  • Media coverage / website / newsletters on the project and clean up reports including photos (before and after) to illustrate the process will be soon put on website.
  • Still discussing with HS2 as this will have an impact on the future use as one of the tips will be effected by this.
  • Next meeting End of March 2017 for further updates.



Arial photo of site (19th Feb 2017)

Download copy of these minutes here (PDF)


Leave a Reply

Fill in your details below or click an icon to log in: Logo

You are commenting using your account. Log Out /  Change )

Facebook photo

You are commenting using your Facebook account. Log Out /  Change )

Connecting to %s